The mass of the Earth of the Earth is approximately 6.0x1024kg while the mass of the Moon is 7.3x1022kg. The distance between both bodies is 3.9x108 m. What is the gravitational force between the Moon and the Earth?
1
Expert's answer
2021-05-11T12:00:52-0400

"F_g=\frac{Gm_1m_2}{r^2} = \frac{6.674\times10^{-11}\times6.0\times10^{24}\times7.3\times10^{22}}{(3.9\times10^8)^2}=1.922\times10^{20} N."